OpenStudy (anonymous):

y varies inversely as twice x. When x = 3, y = 6. Find y when x = 2.

OpenStudy (saifoo.khan):



OpenStudy (anonymous):

soo......?

OpenStudy (saifoo.khan):

soo find the value for k first.

OpenStudy (saifoo.khan):

6 = k/2(3) k = 6*6 k = 36.

OpenStudy (saifoo.khan):

Now, y = 36/x y = 36/2 y = 18 Answ.er

OpenStudy (anonymous):

a. y = 3 b. y = 6 c. y = 12 d. y = 9 I have gotten two answers now, both of which are not a, b, c, or d!